# Stochastic Control Methods for Optimization

### v1

@@ Abstract
In this work, we investigate a stochastic control framework for global optimization over both finite-dimensional Euclidean spaces and the Wasserstein space of probability measures. In the Euclidean setting, the original minimization problem is approximated by a family of regularized stochastic control problems; using dynamic programming, we analyze the associated Hamilton–Jacobi–Bellman equations and obtain tractable representations via the Cole–Hopf transform and the Feynman–Kac formula. For optimization over probability measures, we formulate a regularized mean-field control problem characterized by a master equation, and further approximate it by controlled $N$-particle systems. We establish that, as the regularization parameter tends to zero (and as the particle number tends to infinity for the optimization over probability measures), the value of the control problem converges to the global minimum of the original objective. Building on the resulting probabilistic representations, Monte Carlo-based numerical schemes are proposed and numerical experiments are reported to illustrate the practical performance of the methods and to support the theoretical convergence rates.
@@ 1 Introduction
where $\mathcal{X}$ is a finite-dimensional space $\mathbb{R}^{d}$ or the space of probability measures and the objective function $G:\mathcal{X}\rightarrow\mathbb{R}$ may be non-convex and/or non-differentiable. Optimization problems of the form (1.1) arise in a wide range of applications, including operational research, machine learning, artificial intelligence, finance, engineering design, economics, and operations research. Traditional optimization methods, such as gradient descent and Newton’s method, often struggle with non-convex or non-differentiable functions due to the presence of multiple local minima and the lack of gradient information; see [22, 60]. To address these challenges, we propose and study stochastic control methods to solve the optimization problem (1.1).
@@ 1.1 Optimization over finite-dimensional spaces: $\mathcal{X}=\mathbb{R}^{d}$
$X_{t}^{0,x_{0};\theta}=x_{0}+\int_{0}^{t}\theta_{s}ds+W_{t},\quad t\in[0,1],$ (1.3)
where the $\mathbb{R}^{d}$-valued initial value $x_{0}$ and the $d$-dimensional Wiener process $(W_{t})_{t\geq 0}$ are defined on a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ so that $x_{0}$ is $\mathcal{F}_{0}$-measurable and the Wiener process $(W_{t})_{t\geq 0}$, independent of $x_{0}$, is $(\mathcal{F}_{t})_{t\geq 0}$-adapted. The admissible control $\theta$ is an $(\mathcal{F}_{t})_{t\geq 0}$-adapted process, and the set of admissible controls, denoted by $\Theta$, consists of all such control processes on $[0,1]$ that are square-integrable on $\Omega\times[0,1]$
@@
$u(t,x)=\mathbb{E}\left[e^{-\frac{1}{\varepsilon}G(W_{1}-W_{t}+x)}\right]$ (1.10)
@@ 1.1.3 Main result on convergence analysis as $\varepsilon\to 0$
$\left|V_{\varepsilon}(0,x_{0})-G(\xi)\right|\leq C\,\varepsilon\,\ln\left(\frac{1}{\varepsilon}\right),$ (1.15)
@@ 1.2 Optimization over probability measures: $\mathcal{X}=\mathcal{P}_{2}(\mathbb{R}^{d})$
subject to the controlled SDE (1.3), where $\varepsilon\in(0,1)$ is a regularization parameter.
@@ 1.2.1 The master equation and $N$-Particle Approximation
where $x_{0}\in\mathbb{R}^{d}$ and $\{W^{i}\}_{i=1}^{N}$ are independent standard $d$-dimensional Brownian motions. Then this is nothing but a dN{dN}-dimensional version of the control problem (1.4). The value function $v_{\varepsilon}^{N}$ satisfies the HJB equation:
@@ 1.2.2 Main result on convergence analysis as $\varepsilon\to 0$ and $N\to\infty$
$\left|\frac{1}{N}v^{N}_{\varepsilon}(0,\mathbf{x})-\mathcal{V}_{0}\right|\leq\underbrace{\frac{L_{\varepsilon}}{2N}}_{\text{Particle Error}}+\underbrace{C_{2}\cdot\varepsilon\ln(1/\varepsilon)}_{\text{Regularization Error}},$ (1.26)
where $L_{\varepsilon}$ depends on $\varepsilon$ and the regularity of $G$, and $C_{2}$ depends on $d$, the Lipschitz constant of $G$, and the Wasserstein-2 distance between ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) $\delta_{x_{0}}$ and $\mu^{*}$.
@@ 1.3 Literature, contributions, and organization of the paper
Optimization problem (1.1) is a central topic in applied mathematics and arises broadly in machine learning, artificial intelligence, statistics, and operations research; see, e.g., [7]. Given the vast literature on both modeling and algorithms, we restrict attention to a few related directions. For general non-convex objectives, computing a global minimizer is generally NP-hard [48]. Accordingly, a wide range of approximation methods has been developed for Euclidean optimization, including first-order schemes such as stochastic gradient descent and its variants [54, 49, 53, 59], as well as derivative-free (zeroth-order) approaches and metaheuristics, including particle swarm optimization [32, 39], consensus-based optimization [15, 52], and simulated annealing [41, 58]. For a recent overview, we refer to [22, 60]. Optimization over spaces of probability measures (e.g., $\mathcal{P}_{2}(\mathbb{R}^{d})$) differs fundamentally from Euclidean optimization, as the domain is infinite-dimensional and is naturally endowed with a nontrivial geometry, typically the Wasserstein geometry. Representative approaches include Wasserstein gradient-flow methods [2, 9, 20, 35, 47], particle-based methods [17, 20, 50], and conditional-gradient (Frank–Wolfe) methods [37], among many others.
@@
  • For the Euclidean optimization, convergence results are established, demonstrating that values of the control problems converge to the global minimum of the objective functional in the rate of $\varepsilon\ln\left(\frac{1}{\varepsilon}\right)$ as the regularization parameter $\varepsilon$ approaches zero. Although this Euclidean optimization case may be viewed as a special case of the measure optimization, the proofs and techniques are different and presume only Hölder continuity of the objective function $G$.
  • The optimization over probability measures is approximated through mean-field control problems which is further approximated via $N$-particle systems (equivalently, $N$-player potential games). We establish rigorous convergence results, demonstrating that the value of the controlled finite-particle system converges to the global minimum of the objective functional at the rate of $\frac{C_{\varepsilon}}{N}+\varepsilon\ln\left(\frac{1}{\varepsilon}\right)$ as the regularization parameter $\varepsilon$ approaches zero.
  • Numerical algorithms are proposed based on the probabilistic representations of the value functions and optimal controls. Numerical experiments are conducted to validate the theoretical findings and demonstrate the effectiveness of the proposed methods.
@@
The remainder of this paper is structured as follows. Section 2 develops the stochastic control formulation for global optimization in finite-dimensional Euclidean spaces, provides a detailed convergence analysis as the regularization parameter $\varepsilon\to 0$, establishes the well-posedness of the associated control problems, and reports numerical experiments. Section 3 extends the proposed framework to optimization over the Wasserstein space of probability measures, presenting the mean-field control formulation, deriving the corresponding convergence results (including the $N$-particle approximation), and illustrating the methodology through numerical studies. Finally, the proof of Theorem 3.3 is provided in Appendix A.
@@ 2.1 A convergence analysis when $\varepsilon\to 0$
    The function $G:\mathbb{R}^{d}\to\mathbb{R}$ is nonnegative and there exist constants $L,p>0$ and $\alpha\in(0,1]$ such that
@@
In the above assumption, the nonnegativity of $G$ may be relaxed by the boundedness from below of $G$ and there is no essential difference. The local Hölder continuity and associated growth continuity allow for cases without differentability and convexity.
@@
For the Brownian bridge $B_{t}$, we have
Then
@@
$\leq 2^{q-1}\left(|x_{0}-\xi|^{q}(1-t)^{q}+(1-t)^{q}\mathbb{E}\left[\left|\int_{0}^{t}\frac{1}{1-s}dW_{s}\right|^{q}\right]\right)$
@@
$\leq C_{q}\left\{E\left[|x_{0}-\xi|^{q}\right](1-t)^{q}+d^{q/2}(1-t)^{q/2}t^{q/2}\right\},$
and thus,
@@
Let Assumption 2.1 hold. Then for any $\xi\in\arg\min_{x\in\mathbb{R}^{d}}G(x)$, there exists an admissible control $\tilde{\theta}^{\varepsilon}\in\Theta$ such that the corresponding state process $X^{0,x_{0};\tilde{\theta}^{\varepsilon}}$ satisfies
@@
Theorem 2.2 indicates that as the regularization parameter $\varepsilon$ approaches zero, the expected value of the objective function evaluated at the state process $X_{1}^{0,x_{0};\tilde{\theta}^{\varepsilon}}$ converges to the global minimum of $G$. The convergence rate is of order $\varepsilon\ln(1/\varepsilon)$, which highlights the trade-off between the regularization strength and the accuracy of the approximation to the original optimization problem. Here, the function $G$ may have multiple minimizers. While the minimum value $G(\xi)$ does not depend on $\xi$, the constant $C$ depends on the choice of the minimizer $\xi$ because we use Brownian bridges in the proof but the choice of $\xi$ is arbitrary within the set of minimizers.
@@
Let $\delta\in(0,1)$ whose value is to be determined. Set $\bar{J}^{\tilde{\theta}}=\mathbb{E}\left[J_{\varepsilon}^{\tilde{\theta}^{\varepsilon}}(0,x_{0})\right]$ with the control process $\tilde{\theta}^{\varepsilon}_{t}=\frac{\xi-B_{t}}{1-t}$ for $t\in[0,1-\delta]$ and $\tilde{\theta}^{\varepsilon}_{t}=0$ for $t\in(1-\delta,1)$. Then we have
@@
and thus
$0$ $\leq\bar{J}^{\tilde{\theta}}-G(\xi)$
$=\mathbb{E}\left[G(X_{1}^{0,x_{0};\tilde{\theta}^{\varepsilon}})-G(\xi)+\frac{\varepsilon}{2}\int_{0}^{1}|\tilde{\theta}^{\varepsilon}_{t}|^{2}dt\right]$
@@
By Assumption 2.1 and Lemma 2.1, we have
$I_{1}$ $\leq L\mathbb{E}\left[(1+|W_{1}-W_{1-\delta}+B_{1-\delta}|^{p}+|B_{1}|^{p})|W_{1}-W_{1-\delta}+B_{1-\delta}-B_{1}|^{\alpha}\right]$
$\leq LC\mathbb{E}\left[\left(1+|W_{1}-W_{1-\delta}|^{p}+|B_{1-\delta}|^{p}+|\xi|^{p}\right)\left(|W_{1}-W_{1-\delta}|^{\alpha}+|B_{1-\delta}-B_{1}|^{\alpha}\right)\right]$
$\leq LC\left(\mathbb{E}\left[1+|W_{1}-W_{1-\delta}|^{\frac{2p}{2-\alpha}}+|B_{1-\delta}|^{\frac{2p}{2-\alpha}}+|\xi|^{\frac{2p}{2-\alpha}}\right]\right)^{\frac{2-\alpha}{2}}\left(\mathbb{E}\left[|W_{1}-W_{1-\delta}|^{2}+|B_{1-\delta}-B_{1}|^{2}\right]\right)^{\frac{\alpha}{2}}$
$\leq LC\left(\mathbb{E}\left[1+|W_{1}-W_{1-\delta}|^{\frac{2(p+1)}{2-\alpha}}+|B_{1-\delta}|^{\frac{2(p+1)}{2-\alpha}}+|\xi|^{\frac{2(p+1)}{2-\alpha}}\right]\right)^{\frac{(2-\alpha)p}{2(p+1)}}\left(d\delta+\mathbb{E}\left[|B_{1-\delta}-B_{1}|^{2}\right]\right)^{\frac{\alpha}{2}}$
@@
where $C>0$ is a constant depending only on $p$ and $\alpha$. For $I_{2}$, basic calculations yield
$I_{2}$ $=\frac{\varepsilon}{2}\mathbb{E}\left[\int_{0}^{1-\delta}\left|\frac{\xi-B_{t}}{1-t}\right|^{2}dt\right]$
$=\frac{\varepsilon}{2}\mathbb{E}\left[\int_{0}^{1-\delta}\left|\frac{(\xi-x_{0})(1-t)-(1-t)\int_{0}^{t}\frac{1}{1-s}dW_{s}}{1-t}\right|^{2}dt\right]$
$=\frac{\varepsilon}{2}\mathbb{E}\left[|x_{0}-\xi|^{2}(1-\delta)+\int_{0}^{1-\delta}\left|\int_{0}^{t}\frac{1}{1-s}dW_{s}\right|^{2}dt\right]$
@@
$=\frac{\varepsilon}{2}\mathbb{E}\left[|x_{0}-\xi|^{2}(1-\delta)+d\left(\ln\frac{1}{\delta}-(1-\delta)\right)\right]$
@@
$0\leq\bar{J}^{\tilde{\theta}^{\varepsilon}}-G(\xi)\leq C\,\varepsilon\,\left\{d\ln d+d\ln\left(\frac{1}{\varepsilon}\right)\right\}$
@@
$G(x)-\min_{x\in\mathbb{R}^{d}}G(x)\geq\nu|x-\xi|^{\eta},\quad\text{for }|x-\xi|\leq R;\quad G(x)-G(\xi)\geq\mathcal{E},\quad\text{for }|x-\xi|>R.$ (2.7)
@@
$\leq\frac{C}{\min\{\nu(\gamma\wedge R)^{\eta},\mathcal{E}\}}\,\varepsilon\,\left\{d\ln d+d\ln\left(\frac{1}{\varepsilon}\right)\right\},$

### v2

@@ Abstract
In this work, we investigate a stochastic control framework for global optimization over both Euclidean spaces and the Wasserstein space of probability measures, where the objective function may be non-convex and/or non-differentiable. In the Euclidean setting, the original minimization problem is approximated by a family of regularized stochastic control problems; using dynamic programming, we analyze the associated Hamilton–Jacobi–Bellman equations and obtain tractable representations via the Cole–Hopf transformation and the Feynman–Kac formula. For optimization over probability measures, we formulate a regularized mean-field control problem characterized by a master equation, and further approximate it by controlled $N$-particle systems. We establish that, as the regularization parameter tends to zero (and as the particle number tends to infinity for the optimization over probability measures), the value of the control problem converges to the global minimum of the original objective. Building on the resulting probabilistic representations, Monte Carlo-based numerical schemes are proposed and numerical experiments are reported to illustrate the effectiveness of the methods and to support the theoretical convergence rates.
@@ 1 Introduction
where $\mathcal{X}$ is a finite-dimensional space $\mathbb{R}^{d}$ or the space of probability measures and the objective function $G:\mathcal{X}\rightarrow\mathbb{R}$ may be non-convex and/or non-differentiable. Optimization problems of the form (1.1) arise in a wide range of applications, including operational research, machine learning, artificial intelligence, finance, engineering design, economics, and operations research. Traditional optimization methods, such as gradient descent and Newton’s method, often struggle with non-convex or non-differentiable functions due to the presence of multiple local minimas and the lack of gradient information; see [22, 59]. To address these challenges, we propose and study stochastic control methods to solve the optimization problem (1.1).
@@ 1.1 Optimization over finite-dimensional spaces: $\mathcal{X}=\mathbb{R}^{d}$
$X_{t}^{0,x_{0};\theta}=x_{0}+\int_{0}^{t}\theta_{s}ds+W_{t},\quad t\in[0,1].$ (1.3)
Here and below, the $\mathbb{R}^{d}$-valued initial value $x_{0}$ and the $d$-dimensional Wiener process $(W_{t})_{t\geq 0}$ are defined on a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ so that the filtration $(\mathcal{F}_{t})_{t\geq 0}$ is generated by $x_{0}$ and $(W_{t})_{t\geq 0}$ and augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. In particular, $x_{0}$ is $\mathcal{F}_{0}$-measurable and independent of the Wiener process $(W_{t})_{t\geq 0}$. The admissible control $\theta$ is an $(\mathcal{F}_{t})_{t\geq 0}$-adapted process, and the set of admissible controls, denoted by $\Theta$, consists of all such control processes on $[0,1]$ that are square-integrable on $\Omega\times[0,1]$
@@
$u(t,x)=\mathbb{E}\left[e^{-\frac{1}{\varepsilon}G(W_{1}-W_{t}+x)}\right]\,$, (1.10)
@@ 1.1.3 Main result on convergence analysis as $\varepsilon\to 0$
$0\leq\mathbb{E}\left[V_{\varepsilon}(0,x_{0})\right]-G(\xi)\leq C\,\varepsilon\,\ln\left(\frac{1}{\varepsilon}\right),$ (1.15)
@@ 1.2 Optimization over probability measures: $\mathcal{X}=\mathcal{P}_{2}(\mathbb{R}^{d})$
subject to the controlled SDE (1.3), where $\varepsilon\in(0,\frac{1}{e})$ is a regularization parameter.
@@ 1.2.1 The master equation and $N$-Particle Approximation
where $x_{0}\in\mathbb{R}^{d}$ and $\{W^{i}\}_{i=1}^{N}$ are independent standard $d$-dimensional Brownian motions and the admissible control set $\Theta$ is consisting of addmissible controls $\theta$ that are adapted to the filtration generated by $\{W^{i}\}_{i=1}^{N}$. Then this is nothing but a dN{dN}-dimensional version of the control problem (1.4). The value function $v_{\varepsilon}^{N}$ satisfies the HJB equation:
@@ 1.2.2 Main result on convergence analysis as $\varepsilon\to 0$ and $N\to\infty$
$\left|\frac{1}{N}v^{N}_{\varepsilon}(0,\mathbf{x})-\mathcal{V}_{0}\right|\leq\underbrace{\frac{L_{\varepsilon}}{2N}}_{\text{Particle Error}}+\underbrace{C_{2}\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right)}_{\text{Regularization Error}},$ (1.26)
where $L_{\varepsilon}$ depends on $\varepsilon$ and the regularity of $G$, and $C_{2}$ depends on $d$, the Lipschitz constant of $G$, and the Wasserstein-2 distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between $\delta_{x_{0}}$ and $\mu^{*}$.
@@ 1.3 Literature, contributions, and organization of the paper
Optimization problem (1.1) is a central topic in applied mathematics and arises broadly in machine learning, artificial intelligence, statistics, and operations research; see, e.g., [7]. Given the vast literature on both modeling and algorithms, we restrict attention to a few related directions. For general non-convex objectives, computing a global minimizer is generally NP-hard [47]. Accordingly, a wide range of approximation methods have been developed for Euclidean optimization, including first-order schemes such as stochastic gradient descent and its variants [53, 48, 52, 58], as well as derivative-free (zeroth-order) approaches and metaheuristics, including particle swarm optimization [31, 38], consensus-based optimization [15, 51], and simulated annealing [40, 57]. For a recent overview, we refer to [22, 59]. Optimization over spaces of probability measures (e.g., $\mathcal{P}_{2}(\mathbb{R}^{d})$) differs fundamentally from Euclidean optimization, as the domain is infinite-dimensional and is naturally endowed with a nontrivial geometry, typically the Wasserstein geometry. Representative approaches include Wasserstein gradient-flow methods [2, 9, 20, 34, 46], particle-based methods [17, 20, 49], and conditional-gradient (Frank–Wolfe) methods [36], among many others.
@@
  • For the Euclidean optimization, convergence results are established, demonstrating that values of the control problems converge to the global minimum of the objective functional in the rate of $\varepsilon\ln\left(\frac{1}{\varepsilon}\right)$ as the regularization parameter $\varepsilon$ approaches zero. Although this Euclidean optimization case may be viewed as a special case of the measure optimization, the proofs and techniques are different, the obtained results lay foundation for the measure optimization, and Euclidean optimization does not necessitate mean-field controls.
  • The optimization over probability measures is approximated through mean-field control problems which are further approximated via $N$-particle systems (equivalently, $N$-player potential games). We establish rigorous convergence results, demonstrating that the value of the controlled finite-particle system converges to the global minimum of the objective functional at the rate of $\frac{C_{\varepsilon}}{N}+\varepsilon\ln\left(\frac{1}{\varepsilon}\right)$ as the regularization parameter $\varepsilon$ approaches zero and the particle number $N$ tends to infinity.
  • Numerical algorithms are proposed based on the probabilistic representations of the value functions and optimal controls. Numerical experiments are conducted to validate the theoretical findings and demonstrate the effectiveness of the proposed methods. To reproduce the results, the code is available at
    https://github.com/Jinnqiu/Stochastic-Control-Methods-for-Optimization.
@@
The remainder of this paper is structured as follows. Section 2 develops the stochastic control formulation for global optimization in Euclidean spaces, provides a detailed convergence analysis as the regularization parameter $\varepsilon\to 0$, establishes the well-posedness of the associated control problems, and reports numerical experiments. Section 3 extends the proposed framework to optimization over the Wasserstein space of probability measures, presenting the mean-field control formulation, deriving the corresponding convergence results (including the $N$-particle approximation), and illustrating the methodology through numerical studies. Finally, the proofs for Proposition 3.2 and Theorem 3.3 are provided in Appendix A and Appendix B, respectively.
@@ 2.1 A convergence analysis when $\varepsilon\to 0$
    For the function $G:\mathbb{R}^{d}\to\mathbb{R}$, there exist constants $L,p>0$ and $\alpha\in(0,1]$ such that
@@
The local Hölder continuity condition, together with the imposed growth control, accommodates objective functions that may be non-differentiable and non-convex.
Recall the the Brownian bridge $B_{t}$ connecting $x_{0}$ and $\xi$ in $\mathbb{R}^{d}$:
@@
Applying Burkholder-Davis-Gundy inequality, we have
@@
$=C_{q}\left\{E\left[|x_{0}-\xi|^{q}\right](1-t)^{q}+d^{q/2}(1-t)^{q/2}t^{q/2}\right\},$
and thus, it follows that
@@
Let Assumption 2.1 hold. Then, for each $\xi\in\arg\min_{x\in\mathbb{R}^{d}}G(x)$, there exists an admissible control $\tilde{\theta}^{\varepsilon}\in\Theta$ such that the corresponding state process $X^{0,x_{0};\tilde{\theta}^{\varepsilon}}$ satisfies
@@
Theorem 2.2 indicates that as the regularization parameter $\varepsilon$ approaches zero, the value of the control problem evaluated at the state process $X_{1}^{0,x_{0};\tilde{\theta}^{\varepsilon}}$ converges to the global minimum of $G$. The convergence rate is at least of order $\varepsilon\ln(1/\varepsilon)$, which highlights the trade-off between the regularization strength and the accuracy of the approximation to the original optimization problem. Here, the function $G$ may have multiple minimizers. While the minimum value $G(\xi)$ does not depend on $\xi$, the constant $C$ depends on the choice of the minimizer $\xi$ because we use Brownian bridges in the proof but the choice of $\xi$ is arbitrary within the set of minimizers.
@@
Let $\delta\in(0,1)$ whose value is to be determined. Set $\bar{J}^{\tilde{\theta}}=\mathbb{E}\left[J_{\varepsilon}^{\tilde{\theta}^{\varepsilon}}(0,x_{0})\right]$ with the control process $\tilde{\theta}^{\varepsilon}_{t}=\frac{\xi-B_{t}}{1-t}$ for $t\in[0,1-\delta]$ and $\tilde{\theta}^{\varepsilon}_{t}=0$ for $t\in(1-\delta,1]$. Then we have
@@
and thus, it holds that
$0\leq\bar{J}^{\tilde{\theta}}-G(\xi)$ $=\mathbb{E}\left[G(X_{1}^{0,x_{0};\tilde{\theta}^{\varepsilon}})-G(\xi)+\frac{\varepsilon}{2}\int_{0}^{1}|\tilde{\theta}^{\varepsilon}_{t}|^{2}dt\right]$
@@
By the local Hölder continuity in Assumption 2.1, we have
$I_{1}$ $\leq L\mathbb{E}\left[(1+|W_{1}-W_{1-\delta}+B_{1-\delta}|^{p}+|B_{1}|^{p})|W_{1}-W_{1-\delta}+B_{1-\delta}-B_{1}|^{\alpha}\right].$
Further, using Jensen’s inequality and Lemma 2.1 gives
$I_{1}$ $\leq LC\left(\mathbb{E}\left[1+|W_{1}-W_{1-\delta}|^{\frac{2p}{2-\alpha}}+|B_{1-\delta}|^{\frac{2p}{2-\alpha}}+|\xi|^{\frac{2p}{2-\alpha}}\right]\right)^{\frac{2-\alpha}{2}}\left(\mathbb{E}\left[|W_{1}-W_{1-\delta}|^{2}+|B_{1-\delta}-B_{1}|^{2}\right]\right)^{\frac{\alpha}{2}}$
@@
where $C>0$ is a constant depending only on $p$ and $\alpha$. For $I_{2}$, basic calculation and Itô isometry yield that
$I_{2}$ $=\frac{\varepsilon}{2}\mathbb{E}\left[\int_{0}^{1-\delta}\left|\frac{(\xi-x_{0})(1-t)-(1-t)\int_{0}^{t}\frac{1}{1-s}dW_{s}}{1-t}\right|^{2}dt\right]$
@@
$0\leq\bar{J}^{\tilde{\theta}^{\varepsilon}}-G(\xi)\leq C\,\varepsilon\,\left\{d\ln d+d\ln\left(\frac{1}{\varepsilon}\right)\right\}\,$,
@@
$G(x)-\min_{x\in\mathbb{R}^{d}}G(x)\geq\nu|x-\xi|^{\eta},\quad\text{for }|x-\xi|\leq R;\quad G(x)-G(\xi)\geq\mathcal{E},\quad\text{for }|x-\xi|>R.$
@@
$\leq\frac{C}{\min\{\nu\gamma^{\eta},\mathcal{E}\}}\,\varepsilon\,\left\{d\ln d+d\ln\left(\frac{1}{\varepsilon}\right)\right\},$